Violet buys pies (good 1) and champagne (good 2) with her income of $100. The price of champagne is p2 = $5 per bottle. Draw her budget constraints in each of the following cases assuming that both goods are divisible. Note that since pies is good 1, you should represent the quantity of pies in the horizontal axis.

(a) Pies cost $5 per pie if she buys between zero and 10 pies, and $4 per pie if she buys more than 10 pies.

(b) Pies cost $5 per pie if she buys between zero and 10 pies; if she buys more than 10 pies, each additional pie or (fraction thereof) is half-price, i.e., p1 = $2.5.
Answer #1

a) There are two kinks, one at 10 pies and another at 11 pies. Before this kink the budget equation is 100 = 5X

+ 5Y and after this kink, the budget line changes to 100 = 4X + 5Y

b) There is a kink at 10 pies. Before this kink the budget equation is 100 = 5X + 5Y and after this kink, the budget

line changes to (100 - 5*10) = 2.5(X - 10) + 5Y or 75 = 2.5X + 5Y